OpenStudy (anonymous):

Use the quadratic formula to find two complex conjugate roots of the equation: (1/2x)^2+2x+9/4=0 x=x ∓yi=_____∓_____i please help!!!

jhonyy9 (jhonyy9):

hope that x is in numerator so than multiply both sides by 4 and simplify so and will get 2x2 +8x+9=0 so x_1,_2=(-8+/- sqrt(64-72))/4 =(-8+/- sqrt(-8))/4 =(-8+/- 2isqrt2)/4 = =(-4+/-isqrt2)/2